### v1

@@ 4.7.2 Intra and inter-device variability
up_down parameter
The up_down parameter is defined for the generalized soft bounds model in the simulation environment of the ‘aihwkit’ as the directional bias between the up and down update size ($\Delta G^{+}$ and $\Delta G^{-}$). In addition, the minimum step in each direction d is described by the following expression [44].
@@
Where $\Delta G^{+}$, $\Delta G^{-}$ define the minimum step size in the up and down direction respectively, and $b_{\rm max}$ and $b_{\rm min}$ represent the upper and lower bounds of the conductance. Therefore, considering an independent definition of each device (i.e. zero d-to-d variability) and a normalized conductance range between -1 and 1, the symmetry point device-level characteristic and the up_down analytical parameter are equivalent.

### v2

@@ 4.7.2 Intra and inter-device variability
$up\_down$ parameter
The $up\_down$ parameter is defined for the generalized soft bounds model in the simulation environment of the ‘aihwkit’ as the directional bias between the up and down update size ($\Delta G^{+}$ and $\Delta G^{-}$). In addition, the minimum step in each direction d is described by the following expression [44].
@@
Where $\Delta G^{+}$, $\Delta G^{-}$ define the minimum step size in the up and down direction respectively, and $b_{\rm max}$ and $b_{\rm min}$ represent the upper and lower bounds of the conductance. Therefore, considering an independent definition of each device (i.e. zero d-to-d variability) and a normalized conductance range between -1 and 1, the symmetry point device-level characteristic and the $up\_down$ analytical parameter are equivalent.